OpenStudy (anonymous):

Imagine that you have a rectangular swimming pool, but you have no idea how deep it is. You know that the pool has a volume of 216 m3. You measured the dimensions of the pool and found that it is 15 m long and 6 m wide. A. Calculate the depth of the pool. Show your work.

OpenStudy (anonymous):

@iGreen @iambatman @ganeshie8

OpenStudy (anonymous):

@sreekar369

OpenStudy (sreekar369):

u said tht it is a rectangular pool so ,its volume=lbh as u given volume=216m3=>lbh=216m3 u have given the length=15m and breadth or width =6m then take depth as height now substitute the values in lbh=216m3 and find height

OpenStudy (sreekar369):

@beacon239

OpenStudy (anonymous):

i still don't understand @sreekar369

OpenStudy (abhisar):

How will you calculate the volume this pool?

OpenStudy (sreekar369):

lbh=216 15x6xh=216 h=216/15x6

OpenStudy (sreekar369):

i think @beacon239 u can solve this

OpenStudy (abhisar):

Volume of the rectangular pool = length X Breadth X depth

OpenStudy (anonymous):

i got 2.4 @sreekar369 @Abhisar

OpenStudy (abhisar):

How?You have been given that length and breadth of the pool is 15m and 6m. And they say that the volume is 216\(\sf m^3\). ok so what we do is suppose that depth of the pool is say d. Now this means that, \(\sf 5\times 6\times d=216\\ \Rightarrow d=?\)

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @beacon239 i got 2.4 @sreekar369 @Abhisar \(\color{blue}{\text{End of Quote}}\) That's correct :-)

OpenStudy (abhisar):

Also post these questions in mathematics section from next time :=D

OpenStudy (sreekar369):

correct

OpenStudy (anonymous):

You the real mvp thank you so much @Abhisar @sreekar369

OpenStudy (sreekar369):

np :-)